OpenStudy (anonymous):

The number of students in an English class at the start of the school year was 32. At the end of the semester, there were 24 students in the class. What was the percent of decrease in the number of students? A. 8% B. 25% C. 33.3% D. 75%

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

32-24=8 8/32=1/4=25%

OpenStudy (anonymous):

(32-24)/(32) *100

OpenStudy (annon):

B

OpenStudy (annon):

because of ^^^ reasons

OpenStudy (anonymous):

pecentage change==(first-last)/(first) *00